OpenStudy (anikate):

can someone please tell me all the transformations? ex: horiz shrink, left, up down?

OpenStudy (usukidoll):

if you have a f(x) = x+5 that means I am moving 5 units up. the + sign without () means up f(x) = x-5 I am moving 5 units down. the - sign without () means down f(x) = (x-5) I am moving 5 units to the right . the - sign with () means right f(x) = (x+5) I am moving 5 units to the left. the + sign with () means left. f(x) = -x a negative sign in front means that I am reflecting the graph I forgot about shrink and stretch... It has something to do with the number

OpenStudy (usukidoll):

that's what my textbook had.. but other websites use different notation... with the same meaning

OpenStudy (anikate):

oh cool thanks!! :) ur a savior

jimthompson5910 (jim_thompson5910):

this page seems like a handy reference http://www.mathsisfun.com/sets/function-transformations.html

OpenStudy (usukidoll):

f(x) +5 move up 5 units f(x) -5 move down 5 units (f(x)-5) move right 5 units (f(x)+5) move left 5 units -f(x) reflection on x -axis f(-x) reflection on y axis f(bx) b = number stretches horizontal 0<b<1 stretches, b>1 shrinks bf(x) b = number, stretch vertical b>1 stretch 0<b<1 shrinks

OpenStudy (usukidoll):

or this page http://www.regentsprep.org/regents/math/algtrig/atp9/funclesson1.htm

OpenStudy (anikate):

sounds good! thanks!